### v2

# Matrix approach to generalized ensemble theory

Abstract
Despite significant advances in stochastic thermodynamics, a universal and rigorous ensemble framework for nonequilibrium steady states remains lacking. Here, we provide a concise framework for generalized ensemble theory through a matrix-based approach. By introducing an observation matrix, we show that any discrete probability distribution can be formulated as a generalized Boltzmann distribution, with observables and their conjugate variables serving as basis vectors and coordinates in a vector space. Within this framework, we identify the minimal sufficient statistics required to infer the Boltzmann distribution. The nonequilibrium thermodynamic relations and fluctuation-dissipation relations naturally emerge from this framework. Our findings provide a new approach to developing generalized ensemble theory for nonequilibrium steady-state systems.
Efforts to develop ensemble theories for nonequilibrium complex systems aim to generalize the Boltzmann distribution beyond Gibbs ensembles [1]. Various generalized ensemble theories have been proposed to handle specific classes of systems, such as Hill’s nanothermodynamics for small systems [2, 3, 4], Edwards’ volume ensemble for granular matter [5, 6, 7], and the generalized Gibbs ensemble for integrable systems [8, 9, 10]. However, these frameworks remain largely model-specific, and there is still no unifying formalism that systematically encompasses these diverse ensemble constructions under a common structure. Moreover, classical assumptions of equilibrium ensemble theory, including ergodicity, detailed balance, and thermal equilibrium, often break down or become ill-defined in non-physical systems (e.g., social systems or neural networks) as well as in nonequilibrium systems [11, 12, 13, 14]. Whether ensemble theory can rigorously describe the probability distributions of such systems remains an open question.
Constructing a generalized ensemble framework for complex systems, especially nonequilibrium steady states (NESS), is crucial for understanding nonequilibrium thermodynamics and response behaviors. Approaches like maximum entropy inference [15] and large deviation theory [16] build Boltzmann distributions by constraining observable averages, where key thermodynamic potentials-entropy and free energy-emerge as rate functions and scaled cumulant generating functions [17, 18]. Despite this progress, extending these frameworks to nonequilibrium systems faces significant challenges: constraints are often empirical with ambiguous observable-selection criteria, and determining rate functions for complex systems remains inherently difficult [19, 20]. Some studies have attempted to map nonequilibrium steady states onto equilibrium behavior by introducing an effective temperature [21, 22, 23, 24]. Nevertheless, a unified theoretical framework that rigorously justifies and systematically establishes such correspondences is still missing. Crucially, a general ensemble formalism for NESS, directly analogous to the Gibbs ensemble in equilibrium statistical mechanics, has not yet been developed.
In this letter, we bridge these gaps by establishing a unified matrix formalism for ensemble theory, where any discrete probability distribution is expressed as a generalized Boltzmann form within a vector space. Different ensemble constructions correspond to different choices of observable bases and their conjugate parameters, allowing diverse systems to be described in a common formalism. Importantly, our framework relies only on the minimal assumption of the existence of a well-defined discrete stationary distribution. This makes it naturally applicable to NESS and non-physical systems, where traditional assumptions—such as ergodicity, thermal equilibrium, or physical replicas—do not hold. The conjugate parameters, constructed via matrix transformations, function as tunable variables analogous to effective temperatures. This framework thereby naturally derives nonequilibrium thermodynamic relations and fluctuation-dissipation relations (FDRs). It thus establishes a new framework for probing the thermodynamics and response behavior of systems far from equilibrium.
Matrix representation—For a discrete system with a steady-state probability distribution, the set of microstates is denoted as $\{\sigma_{1},\sigma_{2},...,\sigma_{N}\}$. The steady-state probability distributions of microstates are represented by the vector $\bm{P}=(p_{1},p_{2},\dots,p_{N})^{\top}$. Each microstate has several observables, and the $i$-th observable is denoted by the observable vector $\bm{a}_{i}=(a_{i1},a_{i2},\dots,a_{iN})$, where $a_{ij}$ is the $i$-th observable for $\sigma_{j}$. Suppose that the system has $N$ linearly independent observation vectors, which can be assembled into a full-rank square matrix $\mathbb{A}$. This observation matrix is thus represented as $\mathbb{A}=(\bm{a}_{1},\bm{a}_{2},\dots,\bm{a}_{N})^{\top}$. Therefore, the product of $\mathbb{A}$ and the probability vector $\bm{P}$ yields the vector of observed averages $\bm{O}=(o_{1},o_{2},\dots,o_{N})^{\top}$, expressed as
$\mathbb{A}\bm{P}=\bm{O},$ (1)
where $o_{i}$ denotes the average value of the $i$-th observable. To impose the normalization condition $\sum_{i=1}^{N}p_{i}=1$, a unit observation vector ($a_{1j}=1$ for all microstates) is assigned in the first row of $\mathbb{A}$, yielding $o_{1}=1$. Hence, $\mathbb{A}$ is an $N$-dimensional full-rank matrix with fixed $\bm{a}_{1}$. $\bm{P}$ can be uniquely determined from the vector of observed averages $\bm{O}$, which is
$\bm{P}=\mathbb{A}^{-1}\bm{O}.$ (2)
By taking the negative natural logarithm of each component of $\bm{P}$, one obtains $\bm{I}=-\ln\bm{P}$, which corresponds to the self-information vector in information theory [25]. By multiplying $\bm{I}$ from the left by $({\mathbb{A}^{\top}})^{-1}$, one obtains a vector $\bm{B}=(\mathbb{A}^{\top})^{-1}\bm{I}$ with entries $(b_{1},b_{2},\dots,b_{N})^{\top}$. Then, the self-information vector can be expressed as
$\bm{I}=-\ln\bm{P}=\mathbb{A}^{\top}\bm{B}.$ (3)
Therefore, the probability of microstate $\sigma_{j}$ is $p_{j}=\exp\left(-\sum^{N}_{i=2}b_{i}a_{ij}\right)/\exp(b_{1})$ due to $a_{1j}=1$. This defines a generalized Boltzmann distribution, with the normalization factor $\exp(b_{1})$ identified as the partition function $\mathcal{Z}$. For $i>1$, each row vector $\bm{a}_{i}$ represents a physically measurable observable, and the corresponding coefficient $b_{i}$ is its thermodynamic conjugate variable. For convenience, we refer to the vector $\bm{B}$ as the Boltzmann vector. This vector carries clear physical meaning: its first component, associated with $\ln\mathcal{Z}$, corresponds to a generalized free energy, while each of the remaining components represents a conjugate variable linked to a specific observable—such as the inverse temperature for energy, the chemical potential for particle number, or external magnetic field for spin magnetization. These quantities serve as control parameters that can be tuned to regulate the system.
Eq. (3) shows that an arbitrary discrete probability distribution can be expressed as a generalized Boltzmann distribution through the observation matrix $\mathbb{A}$. It demonstrates that the Boltzmann distribution is a specific representation of probability distributions and is not exclusively confined to equilibrium systems.
Matrix transformations—Eq. (3) can be expressed as $\bm{I}=\sum_{i=1}^{N}b_{i}\bm{a_{i}}^{\top}$, where the vector set $\{\bm{a_{i}}\}_{i=1}^{N}$ form a basis for an $N$-dimensional vector space, with $\bm{B}=(b_{1},b_{2},\dots,b_{N})^{\top}$ acting as coordinates. Since $\bm{a_{1}}$ is fixed, we define the complementary subspace of $\text{span}(\bm{a_{1}})$ as $\mathcal{V}=\text{span}(\{\bm{a_{i}}\}_{i=2}^{N})$. In $\mathcal{V}$, the vectors $\{\bm{a_{i}}\}_{i=2}^{N}$ constitute a complete basis, and $\{b_{i}\}_{i=2}^{N}$ represent coordinates within this subspace. The parameter $b_{1}$ is not independent but is fixed by the normalization condition, given by $\exp(b_{1})=\sum_{j=1}^{N}\exp(-\sum_{i=2}^{N}b_{i}a_{ij})$. Consequently, $\bm{I}$ is uniquely determined by the basis vectors and coordinates within $\mathcal{V}$. This vector space representation demonstrates that distinct choices of basis vectors in $\mathcal{V}$ lead to different coordinate representations $\{b_{i}\}_{i=2}^{N}$ for a given probability distribution. Crucially, the infinite degrees of freedom in selecting basis vectors for $\mathcal{V}$ imply that a probability distribution admits infinitely many equivalent Boltzmann distribution forms.
Under a change of basis in $\mathcal{V}$, $\mathbb{A}$ is transformed to $\mathbb{T}\mathbb{A}$, where $\mathbb{T}$ is an invertible $N\times N$ matrix. Eq. (1) becomes $\mathbb{T}\mathbb{A}\bm{P}=\mathbb{T}\bm{O}$, indicating that both $\mathbb{A}$ and $\bm{O}$ transform by being left-multiplied by $\mathbb{T}$. To preserve normalization (i.e., the first row of $\mathbb{T}\mathbb{A}$ consists of all ones), $\mathbb{T}$ must satisfy the constraint $T_{1j}=\delta_{1j}$ for all microstates. The generalized Boltzmann distribution then takes the form $-\ln\bm{P}=\mathbb{A}^{\top}\mathbb{T}^{\top}(\mathbb{T}^{\top})^{-1}\bm{B}$, where the Boltzmann vector transforms as $\bm{B}\to(\mathbb{T}^{\top})^{-1}\bm{B}$. This means that when the observations change, the probability distribution remains unchanged, leading to gauge freedom in statistical mechanics. See the Supplemental Material (SM) for details.
Figure 1: The illustration of matrix representation of 3-spin model with binary state (±1$\pm$ 1± 1). Spin microstates are ordered from spin 3 to spin 1. The left column shows observables evaluated for each microstate, such as $s_{2}s_{1}$, which denotes the product of spins 2 and 1. Each configuration includes a complete set of observables ranging from single-spin terms to higher-order products, ultimately forming the Hadamard matrix. The first row is the normalized vector.
An example of physically interpretable observation matrices—Although any full-rank matrix with the fixed first row $\bm{a_{1}}$ can mathematically be used to represent a generalized Boltzmann distribution, the observation matrix must be selected to ensure that it is physically interpretable and measurable. Spin models are widely used in combinatorial optimization [26, 27], neural networks [28, 29, 30], and the modeling of biological [31, 32, 33] and social systems [34]. Despite their nonequilibrium or non-Hamiltonian in nature, these systems are often described using equilibrium distributions of spin models. Why are such distributions effective? We address this question by analyzing the matrix representation of spin-model probabilities.
The Sylvester Hadamard matrix $\mathbb{H}$ [35] provides a natural observation matrix for spin models with $n$ binary variables (shown in Fig. 1). Its first row consists entirely of ones, enforcing normalization, while the remaining rows represent spin products, ranging from single-spin observables to full $n$-spin products (see the SM for details). The Boltzmann distribution based on the Hadamard matrix takes the form $-\ln\bm{P}=\mathbb{H}^{\top}\bm{B}$. For a given state $\sigma_{j}$, the probability is
$p_{j}=\frac{1}{\exp{(b_{1})}}\exp{(-\sum_{i=2}^{2^{n}}b_{i}h_{ij})},$ (4)
where $h_{ij}$ denotes the ($i,j$)-th entry of $\mathbb{H}$. This expression can be compared to the Boltzmann distribution of a spin system at equilibrium. The term $\exp(b_{1})$ plays the role of the partition function, while $h_{ij}$ corresponds to the measurement of the $i$-th observable on microstate $\sigma_{j}$, which can represent single-spin quantities, pairwise spin products, and up to $N$-spin products. For $i>1$, the conjugate variable $b_{i}$ represents the dimensionless interaction strength, defined by $b_{i}=J_{i}/k_{B}T$. Interactions $J_{i}$ associated with single-spin observables correspond to external fields, while those associated with multi-spin products represent multi-body interactions.
As shown in Eq. (4), the probability distribution of any $n$-binary system can be expressed in the form of a generalized Boltzmann distribution. This distribution structure directly corresponds to an equilibrium spin model with multi-order interactions. This explains why such a representation remains effective even for nonequilibrium and non-Hamiltonian systems. In practice, second-order spin models are often employed to approximate higher-order spin interactions when modeling complex systems [36, 37]. It is important to note, however, that although these nonequilibrium systems are modeled using equilibrium-like spin models, the inferred interactions effectively encode nonequilibrium characteristics rather than equilibrium properties.
Conservation laws, symmetry and ergodicity breaking reduce the rank of $\mathbb{A}$—Physical systems are often constrained by symmetries and conservation laws, which restrict the probability distribution to a lower-dimensional manifold. These features may introduce linear dependencies within the observation matrix $\mathbb{A}$, thereby reducing its effective rank. For conservation laws, if an observable remains constant across all microstates, its corresponding row vector $\bm{a_{cons}}$ in $\mathbb{A}$ becomes linearly dependent on the first row $\bm{a_{1}}$, which encodes the normalization condition. Including such a row reduces the rank of $\mathbb{A}$. Likewise, symmetries such as rotations or reflections may render distinct microstates observationally indistinguishable, resulting in identical columns in $\mathbb{A}$ and further rank reduction.
Breakdown of ergodicity can also lead to a reduction in the effective dimensionality of $\mathbb{A}$. When certain microstates are dynamically inaccessible, the corresponding probabilities vanish, resulting in divergent self-information $-\ln p_{i}\to\infty$. To ensure that the generalized Boltzmann representation remains well-defined, the analysis must be restricted to the accessible subspace, which allows us to remove the columns of $\mathbb{A}$ associated with zero-probability states. This process induces linear dependencies among the remaining observables and thus reduces the rank of $\mathbb{A}$. A minimal full-rank representation could then be constructed to reflect the true support of the distribution.
By reducing a rank-deficient matrix to a minimal full-rank representation, one obtains a more compact description of the system. While the original high-dimensional full-rank matrix remains mathematically valid, the reduced representation is often more efficient and physically meaningful, especially in systems with strong symmetries or conservation constraints.
Observation and Inference—The inverse problem of determining the Boltzmann vector $\bm{B}$ from the vector of observed averages $\bm{O}$ requires solving
$\bm{B}=-{(\mathbb{A}^{\top})}^{-1}\ln(\mathbb{A}^{-1}\bm{O}).$ (5)
However, in practice, the enormous number of microstates makes it intractable to directly measure all averages and perform matrix computations. Instead, one can assume a known reference distribution $\bm{Q}$ with its corresponding Boltzmann vector $\bm{B^{Q}}$ under the observation matrix $\mathbb{A}$. Meanwhile, the target distribution $\bm{P}$ has the Boltzmann parameter vector $\bm{B}$. By subtracting their self-information vectors, one obtains the relation
$-\ln(\bm{P}/\bm{Q})$ $=\mathbb{A}^{\top}(\bm{B}-\bm{B^{Q}})=\mathbb{A}^{\top}\bm{B^{KL}}$ (6a)
$=\underbrace{b_{1}^{KL}\bm{a_{1}}^{\top}}_{\text{Normalization term}}+\underbrace{\sum_{i=2}^{N}b_{i}^{KL}\bm{a_{i}}^{\top}}_{\text{Difference vector }\bm{L}}.$ (6b)
The quantity $-\ln(\bm{P}/\bm{Q})$ represents the difference vector between $\bm{P}$ and $\bm{Q}$ in the self-information space, and its negative inner product with $\bm{P}$ gives the Kullback-Leibler (KL) divergence $D_{KL}(\bm{P}||\bm{Q})$. The term $\bm{B^{KL}}$ denotes their relative coordinates, where the first component is given by $b_{1}^{KL}\coloneqq b_{1}-b_{1}^{Q}=\ln(\mathcal{Z}/\mathcal{Z}^{Q})$, with $\mathcal{Z}^{Q}$ being the partition function of $\bm{Q}$. The remaining components ($\{b_{i}^{KL}\coloneqq b_{i}-b_{i}^{Q}\}_{i=2}^{N}$) represent the coordinate displacements in $\mathcal{V}$, resulting in a difference vector $\bm{L}=\sum_{i=2}^{N}b_{i}^{KL}\bm{a_{i}}^{\top}$ (shown in Fig. 2). This decomposition reveals that the difference vector $-\ln(\bm{P}/\bm{Q})$ separates into two terms: the normalization term and the difference vector $\bm{L}$ in $\mathcal{V}$. We note that the right side of Eq. (6a) plays the same role as the action in dynamical ensembles [38], as well as the negative entropy production in the detailed fluctuation relation of stochastic thermodynamics [39, 40, 41].
For a given matrix $\mathbb{A}$ and $\bm{Q}$, $\bm{B}$ and $\bm{B^{Q}}$ may share the same components at specific indices, implying that certain entries of $\bm{B^{KL}}$ vanish. We define a set $\mathcal{K}=\{i\mid b_{i}^{KL}\neq 0,\,i\neq 1\}$ with $k$ elements, which identifies the non-zero components of $\bm{B^{KL}}$ (excluding the normalization term). The difference vector in the subspace $\mathcal{V}$ then becomes $\bm{L}=\sum_{i\in\mathcal{K}}b_{i}\bm{a_{i}}^{\top}$. The Boltzmann distribution thus takes the form
$\bm{P}$ $=\frac{\bm{Q}\exp{(-\sum_{i\in\mathcal{K}}b_{i}^{KL}\bm{a_{i}}^{\top}})}{\exp{(b_{1}^{KL})}},$ (7)
where $\exp{(b_{1}^{KL})}$ ensures normalization. This modified Boltzmann distribution is the solution of the minimum KL divergence $D_{KL}(\bm{P}||\bm{Q})$ inference (Abbreviated as minKL inference) [42] under the constraints of observed averages $\{o_{i}\}_{i\in\mathcal{K}}$ (see the SM for derivation). Knowledge of $\bm{Q}$ allows the Boltzmann representation of $\bm{P}$ to be fully determined by a small set of observed averages $\{o_{i}\}_{i\in\mathcal{K}}$. Consequently, $\{o_{i}\}_{i\in\mathcal{K}}$ serve as the minimal sufficient statistics [43, 44] for $\bm{P}$, eliminating redundant observables while preserving all critical information about the distribution. When $\bm{Q}$ is the uniform distribution, $\bm{B^{Q}}$ vanishes except for the first component. Eq. (7) simplifies to $p_{j}=\exp{(-\sum_{i\in\mathcal{K}}b_{i}a_{ij}})/\mathcal{Z}$. This corresponds to Jaynes’ maximum entropy framework: the distribution $\bm{P}$ maximizes entropy subject to the constraints of observed averages $\{o_{i}\}_{i\in\mathcal{K}}$, representing a special case of the minKL inference. Although the reference distribution $\bm{Q}$ is mathematically arbitrary, it is typically chosen to minimize the number of nonzero components in the Boltzmann vector by incorporating known symmetries or constraints. For example, a uniform $\bm{Q}$ naturally applies to highly symmetric systems and recovers classical equilibrium ensembles.
Figure 2: For a system with three microstates $\{\sigma_{1},\sigma_{2},\sigma_{3}\}$, assuming that $\bm{a_{1}}$ is orthogonal to the subspace $\mathcal{V}=\text{span}\{\bm{a_{2}},\bm{a_{3}}\}$. $b_{2}$ and $b_{3}$ serve as coordinates on this plane, while $b_{1}$ is determined as a function of $b_{2}$ and $b_{3}$, forming a curved surface. The reference distribution $\bm{Q}$ and the target distribution $\bm{P}$, as labeled on the surface, can be projected onto the plane $\mathcal{V}$. The vector $\bm{L}$, defined as the difference between mapped points, can be expressed in terms of $\bm{a_{2}}$ and $\bm{a_{3}}$.
When observations are insufficient to fully determine the distribution, the minKL inference behaves analogously to Bayesian updating. The reference distribution $\bm{Q}$ plays the role of a prior $\bm{Q_{0}}$, and the observed average values serve to update this prior. The update process takes the form
$-\ln\bm{Q_{1}}$ $=-\ln\bm{Q_{0}}+\mathbb{A}\bm{B^{KL_{1}}},$ (8a)
$\bm{B^{Q_{1}}}$ $=\bm{B^{Q_{0}}}+\bm{B^{KL_{1}}},$ (8b)
where $\bm{B^{KL_{1}}}$ represents the update vector derived from the minKL inference under limited observations. As more observables are incorporated, both the reference distribution $\bm{Q}_{0}$ and its associated $\bm{B_{Q_{0}}}$ are progressively updated to $\bm{Q_{k}}$ and $\bm{B_{Q_{k}}}$, and the approximation to the true distribution $\bm{P}$ becomes increasingly accurate. In the limiting case where a complete set of sufficient statistics is available, the distribution is fully recovered.
Embedding classical ensembles—In classical ensembles, the reference distribution $\bm{Q}$ is typically taken to be uniform, and the target distribution $\bm{P}$ is characterized by the difference vector $\bm{L}=\sum_{i\in\mathcal{K}}b_{i}\bm{a_{i}}^{\top}$, where only a subset of observables carry nonzero conjugate parameters. For example, the canonical ensemble corresponds to a single observable—the Hamiltonian—with $b_{2}=1/k_{B}T$. The grand canonical ensemble includes both the energy $a_{2j}=E(\sigma_{j})$ and the particle number observable $a_{3j}=N(\sigma_{j})$, where $b_{2}=1/k_{B}T$ and $b_{3}=-\mu/k_{B}T$. The microcanonical ensemble corresponds to the limit $\bm{P}=\bm{Q}$, with all $b_{i}=0$ except the normalization term. Hill’s nanothermodynamics [2, 3, 4] adds subdivision potential as additional observables to describe surface effects in finite systems. The generalized Gibbs ensemble [8, 9, 10], relevant for integrable systems, includes all local conserved quantities as sufficient observables. These ensemble theories capture the sufficient statistical observables that govern the probability distribution of the system. However, for systems with long-range interactions or complex constraints, such observables may be insufficient to fully characterize the probabilistic behavior, ultimately leading to the failure of classical ensemble models. Our framework demonstrates that by constructing a sufficient set of observables, a generalized Boltzmann distribution can still capture complex probability distributions. The bijective mapping between $\bm{B}$ and $\bm{I}$ ensures this generalization. Identifying an appropriate sufficient set of observables for specific complex systems remains an important direction for future investigation.
The matrix-based formalism also accommodates Tsallis statistics [45] by replacing the self-information with the q-deformed form [46], $-\ln_{q}\bm{P}=\mathbb{A}^{\top}\bm{B}$. The Tsallis distribution arises when $\bm{B}$ includes a normalized term $\ln_{q}\mathcal{Z}$ and a nonzero coefficient associated with the vector of Hamiltonian. This matrix representation yields a generalized form of Tsallis distribution, structurally analogous to the generalized Boltzmann form, and shows that both distributions can be embedded within a unified matrix framework. More details are shown in the SM.
Thermodynamic relation and FDRs in NESS—The matrix formalism extends the equilibrium ensemble framework to NESS, enabling a unified description of probability distributions that preserves key thermodynamic properties. The thermodynamic relation in NESS is
$S$ $=-\bm{P}^{\top}\ln\bm{P}=\bm{P}^{\top}\mathbb{A}^{\top}\bm{B}=\bm{O}^{\top}\bm{B}$ (9)
$=\ln\mathcal{Z}+\sum^{N}_{i=2}b_{i}o_{i}.$ (10)
This provides a generalized Legendre-type thermodynamic structure for NESS, in direct analogy with equilibrium ensemble theory. The FDRs connect a system’s linear response to its intrinsic fluctuations. Specifically, consider a set of observables whose fluctuations are described by the covariance
$\chi_{ij}=\mathrm{Cov}(\bm{a_{i}},\bm{a_{j}})=\sum_{k=1}^{N}p_{k}\,a_{ik}a_{jk}-o_{i}o_{j},$ (11)
and whose response is quantified by the susceptibility $\partial o_{i}/\partial b_{j}$. Within our framework, the generalized Boltzmann distribution yields the linear relation
$-\partial o_{i}/\partial b_{j}=\chi_{ij}\quad(i>1,j>1),$ (12)
which remains valid even when the system operates far from detailed balance.
For NESS, several studies characterize nonequilibrium system behavior by introducing an effective temperature or modified FDRs [47, 48, 49, 50, 51, 52]. In contrast, the matrix framework reconstructs state-function-like thermodynamic relations for NESS. This breakthrough enables the FDRs to be expressed in a form that is an exact analogue of equilibrium thermodynamics.
[…]
$p_{i}=\frac{\exp\left(b_{i}w_{i-1,i}-b_{i+1}w_{i+1,i}\right)}{\mathcal{Z}},$ (13)
with boundary conditions $w_{01}=w_{N+1,N}=0$. Under detailed balance conditions ($p_{i}w_{ji}=p_{j}w_{ij}$), all net fluxes vanish, reflecting equilibrium. The bijective mapping between $\bm{P}$ and $\bm{O}$ via the full-rank $\mathbb{J}$ ensures equilibrium distributions yield vanishing net fluxes in $\bm{O}$, while NESS exhibit nonzero net fluxes.
The thermodynamic relation is $S=\ln\mathcal{Z}+\sum_{i=2}^{N}b_{i}\langle J_{i}\rangle$, and the FDR is $-\partial\langle J_{i}\rangle/\partial b_{j}=\langle J_{i}J_{j}\rangle-\langle J_{i}\rangle\langle J_{j}\rangle$. Under the substitution $b_{i}=1/t_{i}$ ($i>1$), it becomes $t_{j}^{2}\partial\langle J_{i}\rangle/\partial t_{j}=\langle J_{i}J_{j}\rangle-\langle J_{i}\rangle\langle J_{j}\rangle$, where $t_{j}$ acts as an effective temperature governing edge flux fluctuations. The covariance $\langle J_{i}J_{j}\rangle-\langle J_{i}\rangle\langle J_{j}\rangle$ quantifies fluctuations in the edge fluxes, while the left-hand side represents the response to effective temperature variation.
This framework defines effective temperature for each net flux and establishes nonequilibrium thermodynamic relation involving entropy, free energy, and flux-temperature conjugate pairs. It retains the structural form of equilibrium thermodynamics while extending it to systems far from equilibrium, offering both interpretability and analytical tractability. Recent studies [53, 41, 54, 55, 56, 57, 52] have concentrated on the response behavior of nonequilibrium Markov processes, establishing several equalities and bounds. Notably, Ref. [55, 56, 57, 52] also employ a matrix approach to study response behavior. In [55], They utilize a transition-rate matrix that is structurally similar to the observation matrix but differs only in the placement of the unit vector, and perturb an equation analogous to Eq. (1). In contrast, our approach leverages Eq. (3) to directly establish the link between fluctuations and response, resulting in the fluctuation-dissipation relation. Detailed mathematical connections are derived in the SM.
Conclusion—We have established a generalized ensemble framework that characterizes thermodynamic behavior through matrix representations of probability distributions, requiring neither ergodicity nor equiprobability assumptions. This extends ensemble theory beyond equilibrium to encompass any system admitting a probabilistic description. While we illustrated examples such as the Hadamard matrix for spin systems, the flux matrix for Markov processes, and the Vandermonde matrix for random walks (shown in the SM), developing a systematic approach that integrates system-specific features—such as symmetries, constraints, or environmental couplings—to yield reduced, physically informed matrix representations remains an important open challenge.
S.G. gratefully acknowledges helpful discussions with Hualin Shi and Yanting Wang (ITP, UCAS), and with Xiao Han (Beijing Jiaotong University).
Appendix A Matrix representation of Tsallis distribution
Tsallis entropy [45] is given by
$S_{q}=k\frac{1-\sum_{i=1}^{N}p_{i}^{q}}{q-1}\quad(\sum_{i=1}^{N}p_{i}=1).$ (S22)
We assume $k=1$ and omit it in the following equations. By introducing the $q$-logarithmic and $q$-exponential functions
$\ln_{q}x$ $\equiv\frac{x^{1-q}-1}{1-q}\quad(x>0;\ln_{1}x=\ln x),$ (S23)
$e_{q}^{x}$ $\equiv[1+(1-q)x]^{\frac{1}{1-q}}_{+}\quad(e_{1}^{x}=e^{x};[z]_{+}\equiv\max(z,0)),$ (S24)
the Tsallis entropy can be written as
$S_{q}=-\sum_{i=1}^{N}p_{i}^{q}\ln_{q}p_{i}.$ (S25)
The constraint associated with the Tsallis distribution is $U=\sum_{i=1}^{N}p_{i}^{q}\epsilon_{i}$ (Curado-Tsallis type constraints). We note that this constraint provides an equivalent description of Tsallis statistics using the escort average constraint $U=\sum_{i=1}^{N}p_{i}^{q}\epsilon_{i}/\sum_{i=1}^{N}p_{i}^{q}$. Under this constraint, the maximum Tsallis entropy yields the Tsallis distribution
$p_{i}=\frac{1}{\mathcal{Z}_{q}}(1-(1-q)\beta\epsilon_{i})^{1/(1-q)}=\frac{1}{\mathcal{Z}_{q}}e_{q}^{-\beta\epsilon_{i}},$ (S26)
where $\mathcal{Z}_{q}$ is the partition function of the Tsallis distribution [46].
Following the same procedure in the main article, the generalized constraints can be written as
$\mathbb{A}\mathbb{T}_{q}\bm{P}=\bm{O}.$ (S27)
The matrix $\mathbb{T}_{q}$ is the transfer matrix which maps each $p_{i}$ to $p_{i}^{q}$ and is defined as $\mathbb{T}_{q}\equiv\bm{P^{q}}\bm{P}^{\top}(\bm{P}\bm{P}^{\top})^{-1}$. The $i$-th entry of vector $\bm{P^{q}}$ is $p_{i}^{q}$. Then we have
$\mathbb{A}\bm{P^{q}}$ $=\bm{O},$ (S28)
$-\ln_{q}\bm{P}$ $=\mathbb{A}^{\top}\bm{B}.$ (S29)
If the vector $\bm{B}$ contains the normalized term $\ln_{q}\mathcal{Z}$ and a parameter $b_{2}$, with all other components set to zero, and the second row of $\mathbb{A}$ encodes the Hamiltonian. The Tsallis distribution naturally emerges in this framework. The Tsallis entropy is $S_{q}=-{\bm{P^{q}}}^{\top}\ln_{q}\bm{P}={\bm{P^{q}}}^{\top}\mathbb{A}^{\top}\bm{B}=\bm{O}^{\top}\bm{B}$. The matrix representation of the Tsallis distribution naturally extends to a generalized Tsallis distribution, analogous to the generalization of the Boltzmann distribution. Therefore, within this unified matrix framework, any probability distribution can, in principle, be represented in either a generalized Boltzmann form or a generalized Tsallis form.
Appendix B Derivation of nonequilibrium response
Since $\mathbb{A}\bm{P}=\bm{O}$, the response of the probability vector with respect to the control parameter $\eta$ is
$\partial_{\eta}\mathbb{A}\cdot\bm{P}+\mathbb{A}\cdot\partial_{\eta}\bm{P}=\partial_{\eta}\bm{O},$ (S30)
$\partial_{\eta}\bm{P}=-\mathbb{A}^{-1}\cdot\partial_{\eta}\mathbb{A}\cdot\bm{P}+\mathbb{A}^{-1}\cdot\partial_{\eta}\bm{O}.$ (S31)
A special choice of $\mathbb{A}$ is the matrix $\mathbb{K}_{n}$ in [55]. The only difference lies in the position of the unit vector row. This is convenient for computing the response, since the term $\partial_{\eta}\bm{O}$ vanishes. By setting $\mathbb{A}$ to be the matrix $\mathbb{K}_{n}$, the above response equation can be directly reduced to the main result (Eq. (6)) in [55].
In the main article, we use $-\ln\bm{P}=\mathbb{A}^{\top}\bm{B}$ to derive a new formula for the response behavior. The response to variations in $b_{i}$ leads to the fluctuation-dissipation relations. If the control parameter is $\eta$, we can express the change in the observation average $o_{i}$ with respect to $\eta$ as
$\partial_{\eta}o_{i}$ $=\sum_{j=2}^{N}\frac{\partial o_{i}}{\partial b_{j}}\frac{\partial b_{j}}{\partial\eta}=-\sum_{j=2}^{N}\chi_{ij}\frac{\partial b_{j}}{\partial\eta}.$ (S32)
Appendix C Vandermonde matrix
The Vandermonde matrix $\mathbb{V}$ is commonly used in polynomial interpolation, where its non-zero determinant ensures the uniqueness of the interpolating polynomial [60]. The entries of $\mathbb{V}$ are defined as $V_{ij}=x_{i}^{j-1}$ with distinct $x_{i}$, ensuring that the first column of $\mathbb{V}$ is a vector of ones and the $N$-dimensional $\mathbb{V}$ is a full-rank square matrix. Consequently, the transpose of $\mathbb{V}$ can be employed as an observation matrix by assigning an observation value $x_{i}$ to each microscopic state, with different observations corresponding to different powers of these observation values. Notably, the vector $\bm{O}$ contains the moments of $x_{i}$ rather than $x_{i}$ itself, indicating that the observation reflects the macroscopic properties. Consequently, $\mathbb{V}^{\top}{\bm{P}}=\bm{O}$ and $-\ln\bm{P}=\mathbb{V}\bm{B}$ can be derived, and the probability distribution of state $\sigma_{i}$ is given by
$p_{i}=\frac{\exp(-\sum^{N}_{j=2}b_{j}x_{i}^{j-1})}{\mathcal{Z}}.$ (S33)
[…]